I want to fill my dogs play ball with foam it is 600mm in dia the foam expands by 20 to 1 so how many litres of foam would i need to fill it.
Many thanks John.

Answer by rothauserc(4718)

the volume of a sphere is = 4/3 * pi * radus^3
the volume of the sphere is 4/3 * 22/7 * 300^3
volume of sphere is = 1.13142857 * 10^8 cubic mm's
the foam expands by 20 to 1 so we divide by 20 to find how much cubic mm's we need to fill the ball
1.13142857 * 10^8 / 20 = 5.657143 * 10^6 cubic mm's of foam
now 1 liter = 1 * 10^6 cubic mm's
so the answer is 5.7 liters